OpenStudy (anonymous):

@jim_thompson5910 Last one! :)

jimthompson5910 (jim_thompson5910):

range = set of all possible outputs = set of all possible y values

OpenStudy (anonymous):

My bad, wrong question!

jimthompson5910 (jim_thompson5910):

(f-g)(5) = f(5) - g(5) = ???

OpenStudy (anonymous):

I don't understand what is "f" and what is "g" here.. I know you are supposed to add somewhere, I believe.

jimthompson5910 (jim_thompson5910):

f(5) is the y value when x = 5 for the f(x) function

jimthompson5910 (jim_thompson5910):

also, f(5) is the value of f(x) when x = 5

jimthompson5910 (jim_thompson5910):

f alone is not a variable it's a function

OpenStudy (anonymous):

@lilia222 check the charts and on the top left corner of the chart there is f(x) and g(x) and how jim told you f(5) - g(5) = ?

jimthompson5910 (jim_thompson5910):

to find f(5), you look in the f(x) table locate the row that has x = 5 the value in the column right next to it is the value of f(5)

OpenStudy (imstuck):

You need to find function equations for these to make it easier to understand, lilia222?

OpenStudy (imstuck):

Maybe I can help by helping you find the equations f(x) and g(x)...it would model more closely, then, what you have been doing thus far...

OpenStudy (anonymous):

woah.. okay, I don't understand because in this problem they don't tell you f(x)= or g(x) = . All they tell you is x=5 , so I'm a bit lost...

jimthompson5910 (jim_thompson5910):

See attached and it may help

OpenStudy (anonymous):

f(x) is just another of saying the following equation. So f(x) in this case is a function that when x=-5 and the results is -11. Basically saying for some undescribed equation f that has variable x, when x =-5 the equation results in -11.

OpenStudy (anonymous):

145-65=80

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

what is 29 - 13 =?

OpenStudy (anonymous):

@lilia222 Does that help? The same goes for the g function. So when you see (f-g)(5). It is saying look at function f(5) and subtract function g(5) from it.

OpenStudy (anonymous):

you dont have to multiply

jimthompson5910 (jim_thompson5910):

you don't multiply 5 and 29 to get 145 you simply use 29 as the value of f(5) ie, f(5) = 29

OpenStudy (imstuck):

lilia222...I CAN HELP YOU FIND THE FUNCTION EQUATIONS IF YOU'D LIKE, MY DEAR!

OpenStudy (imstuck):

This could be so much simpler for you...

OpenStudy (anonymous):

f(x)= f(5) = 29(5) = 145? @IMStuck .. please , I'm semi- lost right now.

jimthompson5910 (jim_thompson5910):

the function notation is throwing you off

jimthompson5910 (jim_thompson5910):

they just want you to subtract the y values of each table when x = 5

OpenStudy (imstuck):

Ok, taking the tables one at a time...the first thing you want to do is to write equations for these functions because that is what you have been doing all along, is working with composite functions, right? Where they already gave you the function.

OpenStudy (anonymous):

Okay and yes.

OpenStudy (imstuck):

Looking at the tables, you can write the equations from the info they gave you. The x is the "x" and the f(x0 is the y. The first thing you have to do is remember the point-slope form of an equation, which is \[y-_{1}=m(x-x _{1})\]where m is the slope. You can find your slope by using the points on the table. The slope formula is \[\frac{ y _{2} -y _{1}}{ x _{2}-x _{1} }\]Start with the first two y points and the first two x points to find the slope.

OpenStudy (imstuck):

\[\frac{ 1-(-11) }{ -2-(-5) }=\frac{ 1+11 }{ -2+5 }=\frac{ 12 }{ 3}=4\]If you do that with the second and third values of y and x you will also see that the slope is 4. So m = 4 in our point-slope equation. Now pick a point for you y1 and x1. I picked (1, 13) because they are both positive. Fill in your equation now with the slope value of 4 and your point (1, 13):

OpenStudy (anonymous):

@IMStuck where are you trying to get to?

OpenStudy (imstuck):

\[y-13=4(x-1)\]Solving this for y you get y = 4x + 9. Or in terms of f(x), you get f(x) = 4x + 9. That is your f(x) function. Now let's find g(x) the same way, by finding first the slope and then the equation.

OpenStudy (anonymous):

... oh Lord; could math just die right now.. like now now >.<

OpenStudy (imstuck):

The slope of the function g(x) is 2, and picking a point (5, 13), your equation is g(x) = 2x + 3. Now subtract f(x) - g(x) when x is 5.

OpenStudy (imstuck):

\[4(5) +9-(2(5)+3)=29-13=16\]

OpenStudy (imstuck):

There...that's your answer.

OpenStudy (anonymous):

@lilia222 See if this helps. I just made it for you.

OpenStudy (imstuck):

I got to the answer, mrfriedrice, the answer!

OpenStudy (anonymous):

you actually didnt have to do that much work but bravo!!!!

jimthompson5910 (jim_thompson5910):

This is becoming a lot more complicated than it has to be. when x = 5, y = 29 for the first table. So f(5) = 29 when x = 5, y = 13 for the second table. So g(5) = 13 Now subtract the two y values: f(5) - g(5) = 29 - 13 = 16 So, (f-g)(5) = 16

jimthompson5910 (jim_thompson5910):

That's all you have to do. Look at the table and find the outputs of the functions when the inputs are x = 5. Then subtract those outputs (f - g)

jimthompson5910 (jim_thompson5910):

The attachment shows what I mean (the first part)

OpenStudy (imstuck):

jim_thompson5910, you are brilliant, sir. However, I think lilia222 was so stuck on the composites the way she was used to seeing them, in equation form, that your simplified result wasn't clicking. Yours is so much simpler, but not if it's not how she learned it.

jimthompson5910 (jim_thompson5910):

true it's helpful to know how to find the functions, but it's not necessary for this problem. Sometimes it's handy to know how to do something quickly and efficient (eg: if you're trying to make a quick decision on something)

OpenStudy (anonymous):

You are ALL uniquely brilliant in your OWN way; I applaud, thank, and appreciate ALL of you for your eagerness to help and dedicated, hard work!! Genuinely, thank you! :)